A humane society reports that 18% of all dogs were adopted from an animal shelter. Consider random samples of 90 dogs. What is the mean and standard deviation of the sampling distribution of dogs adopted from animal shelters if the humane society is correct?
Answer #1

Solution

Given that,

p =18%= 0.18

1 - p = 1-0.18=0.82

n = 90

of the sampling distribution

= p =0.18

standard deviation of the sampling distribution

=  [p ( 1 - p ) / n] = [(0.18*0.82) /90 ] = 0.040

YES CORRECT